OpenStudy (anonymous):

write the equation for a line with a slope of 2/3 that passes through the point (1, -5)

OpenStudy (anonymous):

Hint: Standard equation of a straight line in point-slope form, for slope m and point (p,q) is: \[y-q=m(x-p)\] So plugin your values..

OpenStudy (anonymous):

y-(-5)=2/3(x-1)

OpenStudy (zale101):

Use the point-slope form of an equation of a line: \[y-y_1=m(x-x_1)\] (x1,y1 are the given point and m is the given slope. all u need to do, is plugin your numbers

OpenStudy (anonymous):

Wait I still don't get it. What do I plug in if I only have one set of numbers?

OpenStudy (anonymous):

here let me give you a much simpler explanation. For these types of equations, you need to remember the simple formula for all simple linear equations, y=mx+b. X is the first number in what is the ordered pair that you are given(in this case 1) and y is the second number in that ordered pair(-5) the slope is the m(in this case 2/3.) All you have to do is plug in and solve for b so in this case -5=(2/3)(1)+b. So you simplify that down to -5=2/3+b and you subtract 2/3 from -5 to give you an answer of -5 and 2/3=b, so your final answer is y=2/3x-5(and)2/3 :)